OpenStudy (anonymous):

How can polynomials be classified based on the number of terms and on the degree of the terms?

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@FaiqRaees

OpenStudy (compassionate):

Well, We can separate them into levels. power of 2, binomial power of 3, trinomial power of 4 or greater, polynomial e.g., x^2 - 3 is a binomial While x^3 + x^2 - 3 is a trinomial. We classify them by their largest power.

OpenStudy (compassionate):

Degree is different. We just add the powers. e.g., x^2 + y^2 = 7. the degree is 7. A monomial has one term. So, for exaple, 11x binomial has two, tri has three, and poly is greater and so on.

Directrix (directrix):

Classification by Number of Terms

Directrix (directrix):

Another study chart

Directrix (directrix):

Is this correct: x^2 + y^2 = 7. the degree is 7. @Compassionate

OpenStudy (compassionate):

Oops! gosh, there i go with my typing again ;P x^2 + y^2 = degree of 4.

Directrix (directrix):

A monomial such as x^2 *y^2 has degree 4. But, not the binomial x^2 + y^2.